# Static Spherical Vacuum Solution to Bumblebee Gravity with Time-like VEVs

### v3

@@ Abstract
The static spherical vacuum solution in a bumblebee gravity model where the bumblebee field $B_{\mu}$ has a time-like vacuum expectation value $b_{\mu}$ is studied. We show that in general curved space-time solutions are not allowed and only the Minkowski space-time exists. However, it is surprising that non-trivial solutions can be obtained so long as a unique condition for the vacuum expectation $b^{2}\equiv-b^{\mu}b_{\mu}=2/\kappa$, where $\kappa=8\pi G$, is satisfied. We argue that naturally these solutions are not stable since quantum corrections would invalidate the likely numerical coincidence, unless there are some unknown fine-tuning mechanisms preventing any deviation from this condition. Nevertheless, some properties of these novel but peculiar solutions are discussed, and we show that the extremal Reissner-Nordström solution is a limit of one of our solutions.
I Introduction
@@
In recent decades, increasing interest has been directed toward possible violations of Lorentz symmetry, driven by attempts to formulate a consistent theory of quantum gravity and to understand potential deviations from GR at high energies. Among various approaches, effective field theories incorporating spontaneous Lorentz violation have proven especially fruitful. In this context, Bumblebee gravity has emerged as a minimal yet nontrivial extension of GR, wherein a vector field acquires a nonzero vacuum expectation value (VEV), leading to spontaneous breaking of local Lorentz invariance and diffeomorphism invariance in a controlled manner [1, 2, 3, 4, 5].
The Bumblebee model typically introduces a vector field $B_{\mu,}$ coupled non-minimally to gravity and governed by a potential $V(B_{\mu}B^{\mu}\pm b^{2})$, which determines the vacua of the theory at the classical level. The vacua, determined by the vacuum expectation $\langle B^{\mu}\rangle$ such that $V|_{B_{\mu}=\langle B_{\mu}\rangle}$, do not transform as scalars, thus signaling the spontaneously breaking of Lorentz symmetry. This spontaneous Lorentz symmetry breaking results in modifications to the Einstein field equations, leading to potentially observable signatures in gravitational phenomena. In previous studies, an exact Schwarzschild-like solution in bumblebee gravity is proposed [6]. Also, an exact Kerr-like solution is found [7]. But for both of the solutions, the configurations of the bumblebee field only admit space-like VEVs, i.e., $\langle B^{\mu}B_{\mu}\rangle>0$. Similar results are also obtained and analyzed in detail within metric-affine bumblebee models [8, 9, 10]. The purpose of this work is to investigate the solution to the Bumblebee gravity model with a time-like vacuum expectation value of the bumblebee field.
This literature is organized as follows. In Sec. II, we briefly introduce the action and the equation of motion for bumblebee gravity. In Sec. III, we try to solve the equations of motion in a static spherical field configuration with the bumblebee field obtaining time-like VEVs, and we find that for general VEVs of this type there is no solution for the equations, unless $b=\sqrt{2/\kappa}$. In Sec. IV, we solve the equations when $b=\sqrt{2/\kappa}$, and find two kinds of non-trivial solutions. Sec. V is the discussion and the summary. In this literature, we will adopt the metric signature ($+$)(-$+$)( - + + + ) and also all quantities involved are expressed in natural units $(\hbar=c=1)$.
II Brief Introduction of Bumblebee gravity
@@
$S=$ $\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa}R+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B^{\mu\nu}B_{\mu\nu}-V\right)$
@@
where $b$ is a positive real constant, and $s=\pm 1$ to determine whether the expection value of $B_{\mu}$ is time-like or space-like. In the literature, it is usually assumed that $V$ has (at least one of) its minimum/maximum at 00, thus  [^2] The condition for $V^{\prime\prime}(0)$ plays no role in this work.
$V(0)=0,\ \text{and}\ V^{\prime}(0)=0.$ (3)
@@
Tracing out the equation for $g_{\mu\nu}$, we get $R=-\kappa T^{B}$, and thus the above equation can be expressed as
@@
III Static spherical Vacuum solution for bumblebee gravity with general time-like VEVs
@@
where $\alpha$ and $\beta$ are functions of $r$. We consider a time-like ground $b_{\mu}$ as
@@
So for general cases with $b\neq 0$ and $b\neq\sqrt{2/\kappa}$, we have $\alpha^{\prime}(r)=0$. Substitute it into Eq. (14b), we have $\beta^{\prime}(r)=0$. Thus, there is no non-trivial static spherical solution for bumblebee gravity with general time-like VEVs. This is one of the main results of this work. Next we focus on solutions with the condition $b=\sqrt{2/\kappa}$ satisfied.
IV Solution for bumblebee gravity with VEV $b=\sqrt{2/\kappa}$
Notice that by applying
$\frac{r}{4}\left(\rm{{EQ}_{tt}-\rm{EQ}_{rr}}\right)+\frac{1}{2}\rm{EQ}_{\theta\theta}=0,$ (18)
and replacing $b$ with $\sqrt{2/\kappa}$, we have
$e^{2\beta(r)}-(1+r\alpha^{\prime}(r))^{2}=0.$ (19)
Substitute Eq. (19) into Eq. (14d) and Eq. (15), surprisingly, we can find that all of the equations become the same equation as
@@ IV Solution for bumblebee gravity with VEV $b=\sqrt{2/\kappa}$
IV.1 Solutions with constant $\beta$
@@
where $R_{0}$ is a constant. Substituting it into Eq. (20), we can obtain three solutions as $A=0$, $A=-1$, and $A=\frac{2}{\ell}-1$. The case $A=0$ is the flat Minkowski space-time. The case $A=-1$ is ruled out because $g_{rr}=(1+r\alpha^{\prime}(r))^{2}=0$. When $A=\frac{2}{\ell}-1$, we can get
@@
which means that $r=0$ is a singularity of this solution  [^3] Indeed the other numerator factor $07\ell^{2}-4\ell+8=$ 0 has no real root, therefore the singularity cannot be cured by carefully choosing the value of $\ell$ besides $\ell=2$.. So this solution admits a naked singularity located at $r=0$. The Einstein tensor for the solution Eq. (23) is
$G^{\mu}_{\nu}=\mathrm{diag}\left((\ell^{2}-4)A,-(\ell-2)^{2}A,(\ell-2)^{2}A,(\ell-2)^{2}A\right),$ (25)
@@
IV.2 Solutions with non-constant $\beta$
@@
and substitute it into Eq. (20), we get the equation for $\beta(r)$ as
@@
$g_{\mu\nu}=\mathrm{diag}\left(-\left(1-\frac{R_{0}}{r}\right)^{2},\left(1-\frac{R_{0}}{r}\right)^{-2},r^{2},r^{2}\sin^{2}\theta\right).$ (34)
Actually, this solution coincides with the extremal Reissner–Nordström black hole solution, while the solution for $B_{\mu}$ is not the same: in the Reissner-Nordström solution, $B_{\mu}=(Q/r,0,0,0)$ which obeys $B_{\mu}(r)=0\ \text{for }r\to\infty$; in our case, however, the temporal component of $B_{\mu}$ at the infinity is a non-zero constant $\sqrt{2/\kappa}$. Nevertheless, the difference is merely a consequence of choosing different boundary conditions for $B_{\mu}$.
@@
where $C$ is a constant. In the coordinate $(t,u,\theta,\phi)$, the $uu-$component of the metric is
@@
Now we try to find a new coordinate system $(t,\rho,\theta,\phi)$, satisfying (1) $g_{tt}\cdot g_{\rho\rho}=-1$, (2) $s\rightarrow r$ when $r\rightarrow\infty$. Condition (1) implies that
@@
We finally arrive at the result that, in the coordinate system $(t,\rho,\theta,\phi)$, the metric can be expressed as
@@
$R(\rho)$ $=\left(1-\frac{R_{s}}{\rho}\right)^{\ell/2}\rho,$
@@
$=\frac{(\ell-2)^{2}R_{s}^{2}}{4(\rho-R_{s})^{2\ell}\rho^{8-2\ell}}$
×$\phantom{=}\times\left(48\rho^{2}+32(\ell-3)R_{s}\rho+(56-36\ell+7\ell^{2})R_{s}^{2}\right),$ (45)
so when $\ell>0$ and $\ell\neq 2$, $\rho=R_{s}$ is its singularity, and when $\ell<4$ and $\ell\neq 2$, $\rho=0$ is its singularity. When $\ell=2$, this solution degenerates into the flat space-time, just the same as the case discussed in Sec. IV.1. When $\ell\rightarrow 0$, we return to the extremal Reissner–Nordström black hole solution discussed above.
@@
$G^{\mu}_{\nu}=\mathrm{diag}\left((\ell^{2}-4)A,-(\ell-2)^{2}A,(\ell-2)^{2}A,(\ell-2)^{2}A\right),$ (46)
@@
Still, when $0\leq\ell\leq 2$, the metric satisfies the DEC, WEC, NEC, NDEC, and SEC. It is noteworthy that this is also the sufficient condition for the ADM mass Eq. (44) to be non-negative.
@@
From the solutions Eq. (23) and Eq. (41), we know that there is no Schwarzschild-like solution when the bumblebee field obtains a time-like VEV as $b=\sqrt{2/\kappa}$.
V Discussion and Summary
V.1 The stability of the solutions with VEV $b=\sqrt{2/\kappa}$
@@
Here we argue that the solutions with the VEV $b=\sqrt{2/\kappa}$ discussed in Sec. IV are generally unstable, and thus complete one of our main results: there is no non-trivial static spherical solution for bumblebee gravity with time-like VEVs.
@@
First, there is a necessary condition for the stability of the bumblebee model: the Hamiltonian is bounded from below [11], and thus the functional form of the potential is constrained. To our knowledge, stable models are only known for $V\propto X\equiv B_{\mu}B^{\mu}+sb^{2}$ and $V\propto M(n,2,X/\mu^{2})-1$ ($n\geq 3$ for time-like VEVs), where $\mu^{2}$ is an energy scale of the theory and $M(\alpha,\beta,z)$ is the confluent hypergeometric (Kummer’s) function [12, 13]. However both types of potentials only statisfy the condition $V(0)$ while violate $V^{\prime}(0)=0$, for which one cannot solve the equations of motions by assuming a fixed vacuum expectation value of $B_{\mu}$. On the other hand, the conditions in Eq. (3) can most easily be realized by polynomial functions $V\propto X^{n},\ n>1$, and it is shown that in general there is no lower bound on the Hamiltonian [11, 14, 15] for such kind of potentials. This observation then implies the instability of the solutions with $b=\sqrt{2/\kappa}$.
@@
Second, the VEV condition $b=\sqrt{2/\kappa}$ is such a strict requirement, that any deviation, regardless of how small it is, will ultimately cause the non-trivial solutions degenerate into the Minkowski space-time. This is evident if we observe that Eq. (17) forces $\alpha^{\prime}(r)$ to be zero once $\kappa b^{2}\neq 2$ except at $r=0$. Although the seemingly numerical coincidence is maintained once we set $b=\sqrt{2/\kappa}$ at the classical level, quantum corrections would unavoidably shift the VEV of the bumblebee field $B_{\mu}$, and lead the space-time solutions to decay into the flat space-time. If we focus on the bumblebee field sector, similar to the situation in quantum field theories, loop corrections in general could modify the form of the potential $V\to V_{eff}$, for which the location of the minimum can be different from 00 as in Eq. (3). An immediate consequence is that the VEV gets modified correspondingly $b\to b_{eff}\neq\sqrt{2/\kappa}$. Put differently, the non-trivial solutions with $b=\sqrt{2/\kappa}$ are not stable once quantum fluctuations are taken into account. Of course, this argument fails if, because of some unknown mechanisms of the theory of QG, $b$ is fine-tuned to be exactly $\sqrt{2/\kappa}$, which holds to any order of quantum corrections; or, the gravity coupling constant $\kappa$ would also change by QG corrections such that $b_{eff}=\sqrt{2/\kappa_{eff}}$ throughout. However, even without full understanding of the QG theories, we believe that these mechanisms preventing the non-trivial solutions from decay are quite impossible.
@@
V.2 Violation of the weak cosmic censorship conjecture

### v5

@@ Abstract
The static spherical vacuum solution in a bumblebee gravity model where the bumblebee field $B_{\mu}$ has a one-component time-like vacuum expectation value $b_{\mu}$ is studied. We show that in general curved space-time solutions are not allowed and only the Minkowski space-time exists. However, it is surprising that two non-trivial solutions can be obtained so long as a unique condition for the vacuum expectation $b^{2}\equiv-b^{\mu}b_{\mu}=2/\kappa$, where $\kappa=8\pi G$, is satisfied. One of the solutions contains a naked singularity, while the other exhibits features analogous to a confining potential. We argue that naturally these solutions are not stable since quantum corrections would invalidate the likely numerical coincidence, unless there are some unknown fine-tuning mechanisms preventing any deviation from this condition. Nevertheless, the naked singularities and the photon sphere of these novel but peculiar solutions are discussed, and we show that the extremal Reissner-Nordström solution is a limit of one of our solutions.
keywords:
Bumblebee gravity , time-like VEV , vacuum solution , naked singularity , stability
[…]
[inst1] organization=Department of Physics and Chongqing Key Laboratory for Strongly Coupled Physics, Chongqing University, city=Chongqing 401331, country=P. R. China
1 Introduction
@@
In recent decades, increasing interest has been directed toward possible violations of Lorentz symmetry, driven by attempts to formulate a consistent theory of quantum gravity and to understand potential deviations from GR at high energies. Among various approaches, effective field theories incorporating spontaneous Lorentz violation have proven especially fruitful. In this context, Bumblebee gravity has emerged as a minimal yet nontrivial extension of GR, wherein a vector field acquires a nonzero vacuum expectation value (VEV), leading to spontaneous breaking of local Lorentz invariance and diffeomorphism invariance in a controlled manner [1, 2, 3, 4, 5]  [^1] An alternative framework to incorporate Lorentz invariance violation in gravity is introduce a background Kalb-Ramond field [6, 7, 8]..
The Bumblebee model typically introduces a vector field $B_{\mu,}$ coupled non-minimally to gravity and governed by a potential $V(B_{\mu}B^{\mu}\pm b^{2})$, which determines the vacua of the theory at the classical level. The vacua, determined by the vacuum expectation $\langle B^{\mu}\rangle$ such that $V|_{B_{\mu}=\langle B_{\mu}\rangle}$, do not transform as scalars, thus signaling the spontaneously breaking of Lorentz symmetry. This spontaneous Lorentz symmetry breaking results in modifications to the Einstein field equations, leading to potentially observable signatures in gravitational phenomena. In previous studies, an exact Schwarzschild-like solution in bumblebee gravity is proposed [9]. Also, an exact Kerr-like solution is found [10]. But for both of the solutions, the configurations of the bumblebee field only admit space-like VEVs, i.e., $\langle B^{\mu}B_{\mu}\rangle>0$. Similar results are also obtained and analyzed in detail within metric-affine bumblebee models [11, 12, 13], models with cosmological constant [14], higher-dimensional dS/AdS spacetimes [15], models with topological defects [16], and models with generalized uncertainty principle modifications [17, 18]. Meanwhile, phenomenology of bumblebee gravity is also investigated [19, 18, 20, 21, 15]. The purpose of this work is to investigate the solution to the Bumblebee gravity model with a time-like vacuum expectation value of the bumblebee field.
This literature is organized as follows. In Sec. 2, we briefly introduce the action and the equation of motion for bumblebee gravity. In Sec. 3, we try to solve the equations of motion in a static spherical field configuration with the bumblebee field obtaining one-component time-like VEVs, and we find that for general VEVs of this type there is no non-trivial solution for the equations, unless $b=\sqrt{2/\kappa}$. In Sec. 4, we solve the equations when $b=\sqrt{2/\kappa}$, and find two kinds of non-trivial solutions. Sec. 5 is the discussion and the summary. In this literature, we will adopt the metric signature (+++)(-+++) and also all quantities involved are expressed in natural units $(\hbar=c=1)$.
2 Brief Introduction of Bumblebee gravity
@@
$S=$ $\int d^{4}x\sqrt{-g}$
$\times$ $\left(\frac{1}{2\kappa}R+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B^{\mu\nu}B_{\mu\nu}-V\right)$
@@
where $b$ is a positive real constant, and $s=\pm 1$ to determine whether the expection value of $B_{\mu}$ is time-like or space-like. In the literature, it is usually assumed that $V$ has (at least one of) its minimum/maximum at 0, thus  [^3] The condition for $V^{\prime\prime}(0)$ plays no role in this work.
[…]
@@
Take the trace of the equation for $g_{\mu\nu}$, we get $R=-\kappa T^{B}$, and thus the above equation can be expressed as
@@
As discussed in Ref [22], the static spherical vacuum solution to Bumblebee Gravity can be classified into two classes. For the static spherical field configuration $B_{\mu}=b_{\mu}=(b_{t}(r),b_{r}(r),0,0)$, the only non-vanishing components of field strength tensor $b_{\mu\nu}=\partial_{\mu}b_{\nu}-\partial_{\nu}b_{\mu}$ are $b_{rt}=-b_{tr}=b_{t}^{\prime}(r).$ Rewriting Eq. (10) as
$\frac{1}{\sqrt{-g}}\partial_{r}(\sqrt{-g}b^{\bar{\mu}r})-\frac{\xi}{\kappa}R^{\bar{\mu}\bar{\mu}}b_{\bar{\mu}}=0,$ (11)
where there is no summation for the index $\bar{\mu}=t,r,\theta,\phi$, we can easily see that the $r$- component of Eq. (10) is
$\frac{\xi}{\kappa}R^{rr}b_{r}=0.$ (12)
Eq. (12) shows that the spherically symmetric vacuum solutions in the bumblebee model can be classified into two disjoint classes:
  • 1.
    Class I: $b_{r}\equiv 0$,
  • 2.
    Class II: $R_{rr}\equiv 0$.
For the space-like VEVs, $b_{r}$ cannot be identically zero, so the solutions are classified as belonging to Class II, whereas in the time-like cases, $b_{r}\equiv 0$ can be achieved. In this work, we consider only the Class I solutions. Regarding the Class II solutions in the time-like configuration, one can consult Ref. [23].
@@
3 Static spherical Vacuum solution for bumblebee gravity with general time-like VEVs
@@
where $\alpha$ and $\beta$ are functions of $r$. For the Class I solutions, the $b_{\mu}$ is
@@
So for general cases with $b\neq 0$ and $b\neq\sqrt{2/\kappa}$, we have $\alpha^{\prime}(r)=0$. Substitute it into Eq. (16b), we have $\beta^{\prime}(r)=0$. Thus, there is no non-trivial static spherical Class I solution for bumblebee gravity with general time-like VEVs. This is one of the main results of this work. Next we focus on solutions with the condition $b=\sqrt{2/\kappa}$ satisfied.
4 Solution for bumblebee gravity with VEV $b=\sqrt{2/\kappa}$
In the case of $b=\sqrt{2/\kappa}$, $\ell=\xi b^{2}=\frac{2\xi}{\kappa}$, and Eq. (16d) and Eq. (17) reduce to
$\rm{EQ}_{tt}=$ $(\ell-2)r\alpha^{\prime\prime}(r)+\ell r\alpha^{\prime}(r)^{2}$ (20a)
$-(\ell-2)\alpha^{\prime}(r)\left(r\beta^{\prime}(r)-2\right),$
$\rm{EQ}_{rr}=$ $-(\ell+2)r\alpha^{\prime\prime}(r)-\ell r\alpha^{\prime}(r)^{2}$ (20b)
$\rm{EQ}_{\theta\theta}=$ $-\ell r^{2}\alpha^{\prime\prime}(r)-(2+\ell)r^{2}\alpha^{\prime}(r)^{2}$ (20c)
$+r\alpha^{\prime}(r)(\ell r\beta^{\prime}(r)-2(1+\ell))$
$+2(r\beta^{\prime}(r)+e^{2\beta(r)}-1),$
and
$\rm{EQ}^{B}_{t}=$ $(\ell-2)r\alpha^{\prime\prime}(r)+\ell r\alpha^{\prime}(r)^{2}$ (21)
$-(\ell-2)\alpha^{\prime}(r)(r\beta^{\prime}(r)-2).$
Now we have
$0=$ $\frac{r}{4}\left(\rm{{EQ}_{tt}-\rm{EQ}_{rr}}\right)+\frac{1}{2}\rm{EQ}_{\theta\theta}$ (22)
$=$ $e^{2\beta(r)}-(1+r\alpha^{\prime}(r))^{2},$
which gives
$\beta(r)=\ln\left(1+r\alpha^{\prime}(r)\right),$ (23)
and
$\beta^{\prime}(r)=\frac{\alpha^{\prime}(r)+r\alpha^{\prime\prime}(r)}{1+r\alpha^{\prime}(r)}.$ (24)
Substitute Eq. (23) and Eq. (24) into Eq. (20c) and Eq. (21), surprisingly, we can find that all of the equations become the same equation as
@@ 4 Solution for bumblebee gravity with VEV $b=\sqrt{2/\kappa}$
4.1 Solutions with constant $\beta$
@@
where $R_{0}$ is a constant. Substituting it into Eq. (25), we can obtain three solutions as $A=0$, A=1A=-1, and $A=\frac{2}{\ell}-1$. The case $A=0$ is the flat Minkowski space-time. The case A=1A=-1 is ruled out because $g_{rr}=(1+r\alpha^{\prime}(r))^{2}=0$. When $A=\frac{2}{\ell}-1$, we can get
@@
which means that $r=0$ is a singularity of this solution  [^4] Indeed the other numerator factor $7\ell^{2}-4\ell+8=0$ has no real root, therefore the singularity cannot be cured by carefully choosing the value of $\ell$ besides $\ell=2$.. So this solution admits a naked singularity located at $r=0$. The Einstein tensor for the solution Eq. (28) is
$G^{\mu}_{\nu}=\mathrm{diag}\big($ $(\ell^{2}-4)A,-(\ell-2)^{2}A,$
$(\ell-2)^{2}A,(\ell-2)^{2}A\big),$ (30)
@@
4.2 Solutions with non-constant $\beta$
@@
and substituting it into Eq. (25), we get the equation for $\beta(r)$ as
@@
$g_{\mu\nu}=\mathrm{diag}\Big($- $\big(1-\frac{R_{0}}{r}\big)^{2},\big(1-\frac{R_{0}}{r}\big)^{-2},$
$r^{2},r^{2}\sin^{2}\theta\Big).$ (39)
Actually, this solution coincides with the extremal Reissner–Nordström black hole solution, while the solution for $B_{\mu}$ is not the same: in the Reissner-Nordström solution, $B_{\mu}=(Q/r,0,0,0)$ which obeys $B_{\mu}(r)=0\ \text{for }r\to\infty$; in our case, however, the temporal component of $B_{\mu}$ at the infinity is a non-zero constant $\sqrt{2/\kappa}$. The asymptotic behavior $b_{t}(r)\to\sqrt{2/\kappa}$ follows from the constraint $b_{\mu}b^{\mu}=-b^{2};$ we note that an analogous constant asymptotic value may also appear in the electromagnetic case, as it can be generated by a gauge transformation of $A_{\mu}.$
@@
where $C$ is a constant. In the coordinate $(t,u,\theta,\phi)$, the uuuu-component of the metric is
@@
Now we try to find a new coordinate system $(t,\rho,\theta,\phi)$, satisfying (1) $g_{tt}\cdot g_{\rho\rho}=-1$, (2) $\rho\rightarrow r$ when $r\rightarrow\infty$. Condition (1) implies that
@@
We finally arrive at the result that, in the coordinate system $(t,\rho,\theta,\phi)$, the metric can be expressed as  [^5] This solution is very similar to the Janis-Newman-Winicour metric [24, 25]: $ds^{2}=-(1-R_{s}/r)^{\gamma}dt^{2}+(1-R_{s}/r)^{-\gamma}dr^{2}+(1-R_{s}/r)^{1-\gamma}r^{2}d\Omega^{2}$, with $\gamma=2$-$\ell$, but a major difference is an extra $1-R_{s}/r$ in front of the angular part.
@@
$R(\rho)^{2}$ $=\left(1-\frac{R_{s}}{\rho}\right)^{\ell}\rho^{2},$
@@
$=\frac{(\ell-2)^{2}R_{s}^{2}}{4(\rho-R_{s})^{2\ell}\rho^{8-2\ell}}\times\Big(48\rho^{2}+32(\ell-3)R_{s}\rho$
$\ \ \ \,+(56-36\ell+7\ell^{2})R_{s}^{2}\Big),$ (50)
so when $\ell>0$ and $\ell\neq 2$, $\rho=R_{s}$ is its singularity, and when $\ell<4$ and $\ell\neq 2$, $\rho=0$ is its singularity. When $\ell=2$, this solution degenerates into the flat space-time, just the same as the case discussed in Sec. 4.1. When $\ell\rightarrow 0$, we return to the extremal Reissner–Nordström black hole solution discussed above.
@@
$G^{\mu}_{\nu}=\mathrm{diag}\big($ $(\ell^{2}-4)A,-(\ell-2)^{2}A,$
$(\ell-2)^{2}A,(\ell-2)^{2}A\big),$ (51)
@@
Still, when $0\leq\ell\leq 2$, the metric satisfies the DEC, WEC, NEC, NDEC, and SEC. It is noteworthy that this is also a sufficient condition for the ADM mass Eq. (49) to be non-negative.
@@
From the solutions Eq. (28) and Eq. (46), we know that there is no Schwarzschild-like solution when the bumblebee field obtains a time-like VEV as $b=\sqrt{2/\kappa}$ in the Class I cases.
5 Instability, Naked Singularities, and Photon Spheres
5.1 The stability of the solutions with VEV $b=\sqrt{2/\kappa}$
@@
Here we argue that the solutions with the VEV $b=\sqrt{2/\kappa}$ discussed in Sec. 4 are generally unstable, and thus complete one of our main results: there is no non-trivial Class I static spherical solution for bumblebee gravity with time-like VEVs.
@@
First, there is a necessary condition for the stability of the bumblebee model: the Hamiltonian is bounded from below [26], and thus the functional form of the potential is restricted. To our knowledge, stable models are only known for $V\propto X\equiv B_{\mu}B^{\mu}+sb^{2}$ and $V\propto M(n,2,X/\mu^{2})-1$ ($n\geq 3$ for time-like VEVs), where $\mu^{2}$ is an energy scale of the theory and $M(\alpha,\beta,z)$ is the confluent hypergeometric (Kummer’s) function [27, 28]. However both types of potentials only satisfy the condition $V(0)$ while violating $V^{\prime}(0)=0$, for which one cannot solve the equations of motion by assuming a fixed vacuum expectation value of $B_{\mu}$. On the other hand, the conditions in Eq. (3) can most easily be realized by polynomial functions $V\propto X^{n},\ n>1$, and it is shown that in general there is no lower bound on the Hamiltonian [26, 29, 30] for such kinds of potentials, making the models not stable. Specifically, we consider a quadratic potential and the flat spacetime, $V=\frac{\lambda}{2}(X+sb^{2})^{2}$, which is enough for this work and the discussion of stability. As is shown in Ref. [30], the Hamiltonian reads
$H=\int d^{3}x\left(\frac{1}{2}\left(\Pi^{i}\right)^{2}+\frac{1}{4}B_{ij}B^{ij}+V+2V^{\prime}B_{0}^{2}\right),$ (53)
where $\Pi^{i}$ is the conjugate momentum. Then the Hamiltonian yields an effective potential which can be written as [30]
$\frac{1}{\Lambda^{4}}\mathcal{H}_{V}=\frac{V}{\Lambda^{4}}+2\frac{V^{\prime}}{\Lambda^{2}}f^{2},$ (54)
where $\Lambda$ is some constant of dimensions of mass representing the scale of the theory, and $f=B_{0}/\Lambda$. Explicitly, Eq. (54) reads
$\frac{1}{\Lambda^{4}}\mathcal{H}_{V}=\frac{\lambda}{2}$ $\left(-f^{2}+u^{2}+s\left(\frac{b}{\Lambda}\right)^{2}\right)$
$\times$ $\left(3f^{2}+u^{2}+s\left(\frac{b}{\Lambda}\right)^{2}\right),$ (55)
[…]
@@
Second, the VEV condition $b=\sqrt{2/\kappa}$ is such a strict requirement that any deviation, regardless of how small it is, will ultimately cause the non-trivial solutions to degenerate into the Minkowski space-time. This is evident if we observe that Eq. (19) forces $\alpha^{\prime}(r)$ to be zero once $\kappa b^{2}\neq 2$ except at $r=0$. Although the seemingly numerical coincidence is maintained once we set $b=\sqrt{2/\kappa}$ at the classical level, quantum corrections would unavoidably shift the VEV of the bumblebee field $B_{\mu}$, and lead the space-time solutions to decay into the flat space-time. If we focus on the bumblebee field sector, similar to the situation in quantum field theories, loop corrections in general could modify the form of the potential $V\to V_{eff}$, for which the location of the minimum can be different from 0 as in Eq. (3[31]. An immediate consequence is that the VEV gets modified correspondingly $b\to b_{eff}\neq\sqrt{2/\kappa}$, and hence no non-trivial solutions are admitted. Put differently, the non-trivial solutions with $b=\sqrt{2/\kappa}$ are not stable once quantum fluctuations are taken into account. Consequently, we conclude again that the solutions with non-zero time-like VEVs in this work are unstable. Of course, this argument fails if, because of some unknown mechanisms of the theory of QG, $b$ is fine-tuned to be exactly $\sqrt{2/\kappa}$, which holds to any order of quantum corrections; or, the gravity coupling constant $\kappa$ would also change by QG corrections such that $b_{eff}=\sqrt{2/\kappa_{eff}}$ throughout. However, even without full understanding of the QG theories, we believe that these mechanisms preventing the non-trivial solutions from decay are quite impossible, since such fine-tuning seems to be extremely artificial.
@@
5.2 Violation of the weak cosmic censorship conjecture